# Logarithmic-Depth Quantum Circuits for Hamming Weight Projections

### v2

@@ I Introduction
The ability to perform a coherent Hamming weight measurement has various applications in quantum information science. Realizing a coherent Hamming weight measurement efficiently is useful in information processing tasks like entanglement concentration [18] (see also [19, Chapter 19]), multiparty entanglement concentration [20], coherence concentration [21], multicasting of pure quantum states [22], and compression of pure states [23]. Another application of the coherent Hamming weight measurement is in error detection and correction. In the natural setting of independent and identically distributed (i.i.d.) amplitude damping noise, a coherent Hamming weight measurement can used to detect which error has occurred when using the constant-excitation codes of [24]. For error correction of i.i.d. amplitude damping noise, when using such codes, it is possible to perform the correction step based on the constant-excitation space projected into, by employing a coherent Hamming weight measurement.
@@ II Preliminaries
$=R_{z}\!\left(\frac{-\phi}{2}\right)_{C}\left[\alpha|0\rangle_{C}+\beta|1\rangle_{C}\right]|0\rangle_{S}.$ (29)
Thus, the state of the target qubit is unchanged, and the overall effect is to rotate the control qubit. Thus, a controlled-$R_{z}(\phi)$ gate acting on a state where the target qubit is in the state $|0\rangle$ can be replaced with $R_{z}(\frac{-\phi}{2})$ acting on the control qubit.
@@ III.2 Second Algorithm without Zero Padding
To simplify, we now expand the the two dot products in the previous expression.
@@
$\frac{1}{\sqrt{N+1}}\exp{\!\left(-\frac{\mathrm{i}}{2}\Gamma\right)}\sum_{z,y,x=0}^{N}\exp\!\left(\frac{2\pi\mathrm{i}(x-z)y}{N+1}\right)\\ |z\rangle_{C}P_{x}|\psi\rangle_{S}.$ (62)
@@
$\frac{1}{\sqrt{N+1}}\exp{\!\left(-\frac{\mathrm{i}}{2}\Gamma\right)}\sum_{x=0}^{N}|x\rangle_{C}\otimes P_{x}|\psi\rangle_{S},$ (63)
@@ VI Simulation Results
For the shot-noise simulations, each run used $10,000$ shots. For the noisy simulations, we used the 16-qubit IBM simulator FakeGuadalupeV2. Each run consisted of $10,000$ shots and the median $T_{1}$ and $T_{2}$ values are $7.2\times 10^{-5}$ and $8.8\times 10^{-5}$, respectively. For the quantum device runs, we used the 127-qubit IBM device ibm__\__brisbane. Each run consisted of $10,000$ shots and the median $T_{1}$ and $T_{2}$ values are $2.2\times 10^{-4}$ and $1.3\times 10^{-4}$, respectively.

### v3

@@ I Introduction
The ability to perform a coherent Hamming weight measurement has various applications in quantum information science. Realizing a coherent Hamming weight measurement efficiently is useful in information processing tasks like entanglement concentration [18] (see also [19, Chapter 19]), multiparty entanglement concentration [20], coherence concentration [21], multicasting of pure quantum states [22], and compression of pure states [23]. Another application of the coherent Hamming weight measurement is in error detection and correction. In the natural setting of independent and identically distributed (i.i.d.) amplitude damping noise, a coherent Hamming weight measurement can be used to detect which error has occurred when using the constant-excitation codes of [24]. For error correction of i.i.d. amplitude damping noise, when using such codes, it is possible to perform the correction step based on the constant-excitation space projected into, by employing a coherent Hamming weight measurement.
@@ II Preliminaries
$=e^{\frac{-\mathrm{i}\phi}{4}}\left[e^{\frac{\mathrm{i}\phi}{4}}\alpha|0\rangle_{C}+e^{\frac{-\mathrm{i}\phi}{4}}\beta|1\rangle_{C}\right]|0\rangle_{S}$ (29)
$=e^{\frac{-\mathrm{i}\phi}{4}}R_{z}\!\left(\frac{-\phi}{2}\right)_{C}\left[\alpha|0\rangle_{C}+\beta|1\rangle_{C}\right]|0\rangle_{S}.$ (30)
Due to the leftmost phase $e^{\frac{-\mathrm{i}\phi}{4}}$ being a global phase, it has no physically observable consequence: the state of the target qubit is unchanged, and the overall effect is to rotate the control qubit. Thus, a controlled-$R_{z}(\phi)$ gate acting on a state where the target qubit is in the state $|0\rangle$ can be replaced with $R_{z}(\frac{-\phi}{2})$ acting on the control qubit.
@@ III.2 Second Algorithm without Zero Padding
To simplify, we now expand the two dot products in the previous expression.
@@
$\frac{1}{N+1}\exp{\!\left(-\frac{\mathrm{i}}{2}\Gamma\right)}\sum_{z,y,x=0}^{N}\exp\!\left(\frac{2\pi\mathrm{i}(x-z)y}{N+1}\right)\\ |z\rangle_{C}P_{x}|\psi\rangle_{S}.$ (63)
@@
$\exp{\!\left(-\frac{\mathrm{i}}{2}\Gamma\right)}\sum_{x=0}^{N}|x\rangle_{C}\otimes P_{x}|\psi\rangle_{S},$ (64)
@@ VI Simulation Results
For the shot-noise simulations, each run used $10,000$ shots. For the noisy simulations, we used the 16-qubit IBM simulator FakeGuadalupeV2. Each run consisted of $10,000$ shots and the median $T_{1}$ and $T_{2}$ values were $7.2\times 10^{-5}$ and $8.8\times 10^{-5}$, respectively. For the quantum device runs, we used the 127-qubit IBM device ibm__\__brisbane. Each run consisted of $10,000$ shots and the median $T_{1}$ and $T_{2}$ values are $2.2\times 10^{-4}$ and $1.3\times 10^{-4}$, respectively.